OpenStudy (anonymous):

What is 7.09 × 10^–3 written in standard form? A. 0.00709 B. 0.0709 C. 7090 D. 709,000

OpenStudy (anonymous):

You clearly don't understand these, because I've seen so many of these questions by the same user. Do you need help?

OpenStudy (anonymous):

no just making sure my answer is right :) but thanks

OpenStudy (anonymous):

Okay, what did you get then?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Bro look at the exponent,if is positive you move the decimal point to the rigth and if is negative to the left

OpenStudy (anonymous):

C is wrong

OpenStudy (anonymous):

Yeah it's wrong. If you learn how to do these, it's so simple, you'll be like "Wow."

OpenStudy (anonymous):

is a

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

is A

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

When you have a "-" then you have to go backwards so it would be ... .00709 (A)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ya i just went over this stuff in my college prep class

OpenStudy (mathstudent55):

Look at the pattern below: \(10^3 = 10 \times 10 \times 10 = 1,000\) \(10^2 = 10 \times 10 =100\) \(10^1 = 10\) \(10^0 = 1\) \(10^{-1} = \dfrac{1}{10^1} =\dfrac{1}{10} = 0.1\) \(10^{-2} = \dfrac{1}{10^2} = \dfrac{1}{10 \times 10} = \dfrac{1}{100} = 0.01\) \(10^{-3} = \dfrac{1}{10^3} = \dfrac{1}{1,000} = 0.001\) Now that you see what \(10^{-3}\) means, then \(7.09 × 10^{–3} = 7.09 \times 0.001 = 0.00709\)